A passenger on a train traveling at 95 km/h north walks towards the back of the train at 2.0 km/h relative to the train. The passenger’s velocity relative to the ground is:


1
Expert's answer
2021-10-25T09:56:55-0400

Given:

"v=95\:\rm km/h"

"u=2.0\:\rm km/h"


The passenger’s velocity relative to the ground is:

"v_{\rm rel}=v-u=95-2.0=93\:\rm km/h"